OpenStudy (anonymous):

In parallelogram DEFG, DH = x + 3, HF = 3y, GH = 2x – 5, and HE = 5y + 2. Find the values of x and y. A.) x = 39, y = 14 b.) x = 36, y = 13 c.) x = 14, y = 39 d.) x = 13, y = 36

OpenStudy (anonymous):

@jim_thompson5910 there

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

this is a parallelogram, so the diagonals cut each other in half

jimthompson5910 (jim_thompson5910):

what that means is that DH = HF and EH = HG

jimthompson5910 (jim_thompson5910):

DH = HF x+3 = HF ... plug in DH = x+3 x+3 = 3y ... plug in HF = 3y x = 3y - 3 ... solve for x (by subtracting 3 from both sides)

jimthompson5910 (jim_thompson5910):

EH = HG EH = 2x-5 ... plug in GH = 2x – 5 (note: GH and HG are the same segment) 5y + 2 = 2x-5 ... plug in HE = 5y + 2 (note: EH and HE are the same segment)

jimthompson5910 (jim_thompson5910):

so we have 5y + 2 = 2x-5 if you plug in x = 3y - 3, you get 5y + 2 = 2x-5 5y + 2 = 2(3y - 3)-5 now solve for y and tell me what you get

OpenStudy (anonymous):

@jim_thompson5910 um i dont have a calculator with me right now and im not so good in math

jimthompson5910 (jim_thompson5910):

you don't need a calculator to solve for y

jimthompson5910 (jim_thompson5910):

you just need to isolate it on one side

jimthompson5910 (jim_thompson5910):

are you familiar with solving equations?

OpenStudy (anonymous):

@jim_thompson5910 not really

jimthompson5910 (jim_thompson5910):

here's a page that might be helpful http://www.purplemath.com/modules/solvelin.htm

OpenStudy (anonymous):

@jim_thompson5910 so 5y-2=2(3y+3)5 thats going to be the equation

jimthompson5910 (jim_thompson5910):

no the equation is given above